# Experimental practical quantum tokens with transaction time advantage

### v1

@@ 0.1 The practical quantum token scheme
Although not needed in our experimental implementation, it is useful to mention that the scheme can be straightforwardly extended to allow Alice to report losses to Bob. This requires the following extra steps in the quantum phase for the token generation [14]. $A_{0}$ reports to $B_{0}$ the set $\Lambda$ of indices $k$ of quantum states $\lvert\psi_{k}\rangle$ sent by $B_{0}$ that produce unsuccessful measurements. Let $n=\lvert\Lambda\rvert$. $B_{0}$ does not abort if and only if $n\geq\gamma_{\text{det}}N$, where the threshold $\gamma_{\text{det}}\in(0,1)$ is agreed in advanced by Alice and Bob. The scheme continues as above but with the $n-$bit strings that restrict $\textbf{x},\textbf{x}^{\prime},\textbf{x}^{0},\textbf{x}^{1},\textbf{t},\textbf{u}$ to entries with indices $k\in\Lambda$. $P_{\text{det}}\in(0,1)$ is the probability that a quantum state transmitted by Bob is reported by Alice as being successfully measured. The strategy used by $A_{0}$ to report (un)successful measurements must be chosen carefully to counter multi-photon attacks by $B_{0}$ [14, 15]. The analysis for our implementation reduces straightforwardly to the case $P_{\text{det}}=\gamma_{\text{det}}=1$ and $n=N$.
@@ 0.3 Quantum and comparative advantage
    $B_{0}$ gives a classical password y to $A_{0}$ at $L_{0}$, keeps a copy of y and sends a copy to $B_{1}$; $A_{0}$ keeps a copy of y and sends a copy to $A_{1}$.
@@
    $A_{0}$ obtains $b$, indicating that she wishes to present the token at $L_{b}$. She keeps a copy of $b$ and sends a copy to $A_{1}$ at the time $T_{\text{begin,C}}$.
@@
    At the time $T_{\text{bit}}$, $A_{0}$ indicates to $B_{0}$ that the token will be presented at time $T$ given by (1) at either $L_{0}$ or $L_{1}$.
@@
    $A_{b}$ gives y to $B_{b}$ at $L_{b}$ within the time interval $[T,T+\delta T]$.
@@
    At the time $T+\delta T$, $B_{i}$ sends $r_{i}$ to $B_{i\oplus 1}$, where $r_{i}=1$ ($r_{i}=0$) indicates that $B_{i}$ received (did not receive) a token at $L_{i}$ within the time interval $[T,T+\delta T]$.
@@
    At the time $T_{\text{C,end}}=T+\delta T+\Delta T_{\text{comm}}$, $B_{i}$ validates a token received at $L_{i}$ within $[T,T+\delta T]$ if it is equal to y and if $r_{i\oplus 1}=0$.
@@ 0.4 Experimental quantum and comparative advantage
as mentioned in the main text , since $\Delta T_{\text{proc}}\approx 1.5~{}\mu\text{s}$. We observed the quantum advantage in the fiber optic network within the city of Jinan, Shandong Province, as depicted in Fig. 2(a). The fiber length connecting these locations is $L_{\text{fibre}}=2,766$ m, while the direct free-space distance is $D\approx 426$ m. The token is tested 20 times with randomness in choosing the presentation location $b$, resulting in an average transaction time of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta}T_{\text{tran}}=15.336\pm 0.009~{}\mu\text{s}$ and an average error rate of $6.02\%$. Notably, the quantum advantage in this scenario amounts to $QA=12.324\pm 0.009~{}\mu\text{s}$, shown in Fig. 3(a).
@@
as mentioned in the main text. We realize the comparative advantage in field-deployable fibre between Yiyuan ($36^{\circ}10^{\prime}50.4^{\prime\prime}N,118^{\circ}12^{\prime}10^{\prime\prime}E$) and Mazhan ($36^{\circ}0^{\prime}19^{\prime\prime}N,118^{\circ}42^{\prime}35^{\prime\prime}E$) in Shandong Province of China, which are connected by an optical fibre channel of length $L_{\text{fibre}}=60.54$ km and separated by a physical distance of $D=51.60$ km, as shown in Fig. 2(b). We ran the transactions 20 times. The error rates of all trials are below $\gamma_{err}=9.4\%$, with an average error rate of $6.00\%$. Additionally, the average transaction time is ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta}T_{\text{tran}}=304.202\pm 0.011~{}\mu\text{s}$. Thus, we obtained the comparative advantage $CA=39.798\pm 0.011~{}\mu\text{s}$ shown in Fig. 3(b). See SI for further details.
@@ 1 Summary
As discussed in section 7, our quantum token schemes and the security analysis extend straightforwardly to an arbitrary number of presentation spacetime regions. As we discussed, our experimental setup would guarantee a high degree of security in realistic multi-node scenarios involving global or national networks.
@@
where $P_{\text{bound}}^{k}\leq P_{\text{bound}}<1$ for all $k\in[N]$. This further implies that we can upper bound the right-hand side by a Chernoff bound decreasing exponentially with $N$ if $n<N(1-P_{\text{bound}})$.
This result is quite useful for a great variety of quantum cryptography protocols in which Alice’s cheating probability reduces to winning the described task. In this case, the security proof can be reduced to finding the upper bound $P_{\text{bound}}^{k}$ for the round $k$ conditioned on any quantum inputs $\rho_{j}$ and classical outputs $x_{j}$ for rounds $j\neq k$ and on any extra measurement outcome $o_{\text{extra}}$ obtained by Alice, for all $k\in[N]$. Crucially, we note that the result applies to arbitrary quantum strategies by Alice, including arbitrary joint quantum measurements on the quantum states received in all $N$ rounds.
@@ 2 Estimation of experimental imperfections
where $N_{tu}$ is the total number of pulses prepared by Bob in the state labeled by $t$ in the basis labeled by $u$ that were measured by Alice in the basis labeled by $u$, and where $N_{tu}^{\text{error}}$ is the number of Alice’s incorrect measurement outcomes $t\oplus 1$ obtained from the $N_{tu}$ pulses, for all $t,u\in\{0,1\}$. Given that the frequencies $\bar{E}_{tu}$ are obtained from binomial distributions, their standard deviations are
@@
Table S1: Statistics for the average error rates
@@ 2.5 Derivation of an upper bound on $\mu$ from a lower bound on $\eta_{\text{B}}$
3 Upper bound on the uncertainty angle of the prepared state
In this section, we describe the experimental procedure used to determine an upper bound $\theta$ for the uncertainty angle on the Bloch sphere for Bob’s prepared states, and an upper bound $P_{\theta}$ on the probability that this bound is not satisfied.
As shown in Fig. 1 of the main text, Bob prepared the quantum states using two Pockels cells to modulate the bases and encoded bits. The Pockels cells were driven by quantum random number generators (QRNGs). We label the four target states by ‘0’, ‘1’, ‘+’ and ‘-’. Alice used a half-wave plate (HWP) and a polarizing beam splitter (PBS) to measure the quantum states. We used Alice’s setup to measure the quantum states in one of two bases by setting the HWP at one of two possible angles using a rotation mount. We experimentally estimated $\theta$ and $P_{\theta}$ using this joint setup, with some variations discussed below. Given the experimental setup, we refer to these below as measurements by Alice. Note however that in a real-world implementation, these estimations would be performed by Bob using his own independent measurement setup.
To estimate $\theta$, Alice’s two single-photon detectors were replaced by two power meters, and the intensity of the incoming light pulse was set to the higher value of approximately 18 mW. For each of the four states prepared by Bob, we measured the intensity of light measured by each of the two power meters. We repeated this 1000 times for each of the four states prepared by Bob. If Bob prepared the target states perfectly and the experimental setup was ideal then only one of the two power meters would measure a non-zero value. However, due to imperfections in the preparation procedure and the experimental setup, both power meters measure non-zero values, although one is much smaller than the other.
@@ 3 Upper bound on the uncertainty angle of the prepared state
to six decimal places. The uncertainty angle for Bob’s state preparation would be given by $\theta=\max_{i\in\{0,1,+,-\}}\{\alpha_{i}\}$, giving $\theta={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3.429185^{\circ}}$. However, we need to consider the imperfections of the PBS, HWP, and rotation mount to derive a more accurate value of $\theta$, as discussed below.
3.1 Considering the experimental imperfections of the PBS, HWP, and rotation mount
We now provide a derivation of $\theta$ and $P_{\theta}$ considering the experimental imperfections of the PBS, HWP, and rotation mount using the experimental setup depicted in Fig. S1.
@@ 3.1 Considering the experimental imperfections of the PBS, HWP, and rotation mount
We placed the PBS before the polarizer, as shown in Fig. S1(b). The light after the PBS is supposed to be horizontally(vertically)-polarized. We set the polarizer to angles of 0 and 90 degrees to test the actual polarization. We measured the residual vertical component when the light was supposed to be horizontally-polarized, and vice versa. We repeated this procedure ten times. We obtained a contrast of intensities given by
@@
Now let us consider that the states $\ket{+}$ or $\ket{-}$ are prepared. Thus, the HWP is aimed at $22.5^{\circ}$. In this case, if the HWP were perfect and perfectly aligned, the Bloch vector of the imperfectly prepared state would be rotated to the ‘attempted rotated state’ shown in Fig. S2(b). Due to the imperfections and misalignment of the HWP, the state is in fact rotated to the ‘rotated state’ illustrated in Fig. S2(b). Fig. S2(b) illustrates the worst-case scenario that allows us to derive an upper bound on $\theta$. As in Fig. S2(a), we assume the two Bloch vectors lie on the real plane defined by the PBS axes. Thus, as above, (72) and (73) hold.
@@ 4 Time sequence and transaction times
Table S2 gives details of the obtained time measurements relevant for computing QA in our intracity implementation in Jinan, with laboratories communicated by $L_{\text{fibre}}=2,766$ m of optical fibre and physically separated by 425 m. The average value for the transaction time $\Delta T_{\text{tran}}$ was $15.34~{}\mu s$. The transaction time of the classical cross-checking scheme discussed in the Methods section is given by $\Delta T_{\text{tran,C}}=2\Delta T_{\text{comm}}$ (Eq. (8) in Methods), where $\Delta T_{\text{comm}}=\frac{L_{\text{fibre}}}{c_{\text{fibre}}}$ is the time that it takes to communicate a bit between $L_{0}$ and $L_{1}$ over the optical fiber channel, which took the value $\Delta T_{\text{tran,C}}=27.66~{}\mu s$ in our implementation in Jinan, and where $c_{\text{fibre}}=2\times 10^{8}$ ms-1 is the speed of light through the optical fibre. Thus, we obtained a quantum advantage of $QA=\Delta T_{\text{tran,C}}-\Delta T_{\text{tran}}=27.66~{}\mu\text{s}-15.34~{}\mu\text{s}=12.32~{}\mu\text{s}$ in our intracity implementation in Jinan.
Table S3 gives details of the obtained time measurements relevant for computing CA in our intercity implementation between Yiyuan and Mazhan, with laboratories communicated by 60.54 km field-deployable optical fiber and physically separated by $D=51.60$ km. The average value for the transaction time $\Delta T_{\text{tran}}$ was $304.20~{}\mu$s. The transaction time of the classical cross-checking scheme discussed in the Methods section is given by $\Delta T_{\text{tran,CF}}=\frac{2D}{c}=344~{}\mu$s (Eq. (9) in Methods), where $c=3\times 10^{8}~{}$ms-1 is the speed of light through a vacuum. Thus, we obtained a comparative advantage of $CA=\Delta T_{\text{tran,CF}}-\Delta T_{\text{tran}}=344~{}\mu\text{s}-304.20~{}\mu\text{s}=~{}39.80~{}\mu\text{s}$ in our intercity implementation.
@@ 5 Related work and extensions of our schemes
Bartkiewicz et al. [24] describe an experimental implementation of partial cloning attacks on photon states representing components of quantum money tokens. The tokens were very short-lived, as they did not implement quantum memory. Bozzio et al. [26] demonstrated an on-the-fly version of quantum money tokens using weak coherent states of light. Guan et al. [27] also implemented short-lived quantum money with light pulses, using high-dimensional time-bin qudits. Again, neither of these implementations was integrated with quantum memory, and so the tokens were very short-lived. Jirakova et al. [28] presented proof-of-concept attacks on the implementation reported in Ref. [26].
@@
More directly comparable to our experiment is that of Schiansky et al. [35], who recently reported an experimental demonstration of quantum digital payments using the essential concepts of quantum S-money. Ref. [35] also offers comments on and comparisons with our scheme and a previous partial implementation [14]. Detailed comparisons of the merits of quantum money, token and digital payment schemes are indeed important for the progress of the field, and dialogue is essential. We leave a complete review for future work but comment briefly here on some key points.
While interesting, in its present form the scheme of Ref. [35] does not provide any fundamental advantage over a purely classical scheme. This is because the insecure quantum channel over which the bank (token provider in the notation of Ref. [35]) sends the quantum states that generate the token to the client can be straightforwardly replaced by a secure classical channel using previously distributed keys. This then allows a scheme based on purely classical tokens combined with cross-checking. Note in this context that, as presented, the scheme of Ref. [35] does not satisfy instant validation, as it requires two-way communication between distant locations to complete a transaction (the merchant receiving a token from the client communicates with the bank for payment verification and the bank must communicate back to the merchant to transfer the money). While one could consider an alternative scenario in which merchants are able to validate without reference to the bank, the merchants then require the bank’s token data and effectively become local agents of the bank, which is the scenario discussed in Refs.  [11, 14] and implemented in our experiment.
Note further that the implementation of Ref. [35] required “a few tens of minutes” [35] from Alice’s (the client’s) choice of presentation point to token validation. This should be compared with the $\approx 1.5~{}\mu{\rm s}$ required to validate a presented token in our experiments, the total transaction times of $\approx 15~{}\mu{\rm s}$ and $\approx 30{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}~{}\mu{\rm s}$ for our intra-city and inter-city experiments and also, importantly, with the times required for classical cross-checking schemes to complete a token transaction, which would be $\approx 28~{}\mu{\rm s}$ in our intracity setup using our 2,766 m long optical fibre link and $\approx 344~{}\mu{\rm s}$ in our intercity scenario using ideal light speed communication through free space with nodes separated by 51.651.6 51.651.6 km, and which would be $\approx 7~{}\mu{\rm s}$ for the 641 m long optical fibre link used in the experiment of Ref. [35] – or even shorter if the classical cross-checking scheme used ideal light-speed communication through free space. As we have emphasized, the practical motivation for implementing a quantum token scheme relies on being able to demonstrate advantage compared to the classical cross-checking.
@@
Another issue is that the implementation of Ref. [35] requires the user/client to choose, at the time they receive the quantum token data from the bank, the location/merchant for which the token will be valid. In contrast, our recent [14] and current implementation allow the user flexibility to make this choice shortly before presenting the token for verification. The time between choice and presentation is limited in principle only by causal communication constraints, and in practice we achieved the sub-millisecond transaction times reported above. Our implementation allows the quantum communication between bank and user to take place arbitrarily long (even years) before the user chooses when and where to spend the token. In this respect, it replicates the real-world functionality of money and credit cards, which are typically obtained from banks without any need to commit to using them with a given merchant or at a given time and place.

### v2

@@ 0.1 The practical quantum token scheme
Although not needed in our experimental implementation, it is useful to mention that the scheme can be straightforwardly extended to allow Alice to report losses to Bob. This requires the following extra steps in the quantum phase for the token generation [14]. A0 reports to B0 the set $\Lambda$ of indices $k$ of quantum states $\lvert\psi_{k}\rangle$ sent by B0 that produce unsuccessful measurements. Let $n=\lvert\Lambda\rvert$. B0 does not abort if and only if $n\geq\gamma_{\text{det}}N$, where the threshold $\gamma_{\text{det}}\in(0,1)$ is agreed in advanced by Alice and Bob. The scheme continues as above but with the $n-$bit strings that restrict $\textbf{x},\textbf{x}^{\prime},\textbf{x}^{0},\textbf{x}^{1},\textbf{t},\textbf{u}$ to entries with indices $k\in\Lambda$. $P_{\text{det}}\in(0,1)$ is the probability that a quantum state transmitted by Bob is reported by Alice as being successfully measured. The strategy used by A0 to report (un)successful measurements must be chosen carefully to counter multi-photon attacks by B0 [14, 15]. The analysis for our implementation reduces straightforwardly to the case $P_{\text{det}}=\gamma_{\text{det}}=1$ and $n=N$.
@@ 0.3 Quantum and comparative advantage
    B0 gives a classical password y to A0 at $L_{0}$, keeps a copy of y and sends a copy to B1; A0 keeps a copy of y and sends a copy to A1.
@@
    A0 obtains $b$, indicating that she wishes to present the token at $L_{b}$. She keeps a copy of $b$ and sends a copy to A1 at the time $T_{\text{begin,C}}$.
@@
    At the time $T_{\text{bit}}$, A0 indicates to B0 that the token will be presented at time $T$ given by (1) at either $L_{0}$ or $L_{1}$.
@@
    Ab gives y to Bb at $L_{b}$ within the time interval $[T,T+\delta T]$.
@@
    At the time $T+\delta T$, Bi sends $r_{i}$ to Bi⊕1, where $r_{i}=1$ ($r_{i}=0$) indicates that Bi received (did not receive) a token at $L_{i}$ within the time interval $[T,T+\delta T]$.
@@
    At the time $T_{\text{C,end}}=T+\delta T+\Delta T_{\text{comm}}$, Bi validates a token received at $L_{i}$ within $[T,T+\delta T]$ if it is equal to y and if $r_{i\oplus 1}=0$.
@@ 0.4 Experimental quantum and comparative advantage
as mentioned in the main text , since $\Delta T_{\text{proc}}\approx 1.5~{}\mu\text{s}$. We observed the quantum advantage in the fibre optic network within the city of Jinan, Shandong Province, as depicted in Fig. 2(a). The fibre length connecting these locations is $L_{\text{fibre}}=2.766$ km, while the direct free-space distance is $D\approx 426$ m. The token is tested 20 times with randomness in choosing the presentation location $b$, resulting in an average transaction time of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta}T_{\text{tran}}=15.34\pm 0.01~{}\mu\text{s}$ and an average error rate of $6.02\%$. Notably, the quantum advantage in this scenario amounts to $QA=12.32\pm 0.01~{}\mu\text{s}$, shown in Fig. 3(a).
@@
as mentioned in the main text. We realize the comparative advantage in field-deployable fibre between Yiyuan ($36^{\circ}10^{\prime}50.4^{\prime\prime}N,118^{\circ}12^{\prime}10^{\prime\prime}E$) and Mazhan ($36^{\circ}0^{\prime}19^{\prime\prime}N,118^{\circ}42^{\prime}35^{\prime\prime}E$) in Shandong Province of China, which are connected by an optical fibre channel of length $L_{\text{fibre}}=60.54$ km and separated by a physical distance of $D=51.60$ km, as shown in Fig. 2(b). We ran the transactions 20 times. The error rates of all trials are below $\gamma_{err}=9.4\%$, with an average error rate of $6.00\%$. Additionally, the average transaction time is ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta}T_{\text{tran}}=304.20\pm 0.01~{}\mu\text{s}$. Thus, we obtained the comparative advantage $CA=39.80\pm 0.011~{}\mu\text{s}$ shown in Fig. 3(b). See SI for further details.
@@ 1 Summary
As discussed in section 7, our quantum token schemes and the security analysis extend straightforwardly to an arbitrary number of presentation spacetime regions. As we discuss, our experimental setup would guarantee a high degree of security in realistic multi-node scenarios involving global or national networks.
@@
where $P_{\text{bound}}^{k}\leq P_{\text{bound}}<1$ for all $k\in[N]$. This further implies that we can upper bound the right hand side by a Chernoff bound decreasing exponentially with $N$ if $n<N(1-P_{\text{bound}})$.
This result is quite useful for a great variety of quantum cryptography protocols in which Alice’s cheating probability reduces to wining the described task. In this case the security proof can be reduced to finding the upper bound $P_{\text{bound}}^{k}$ for the round $k$ conditioned on any quantum inputs $\rho_{j}$ and classical outputs $x_{j}$ for rounds $j\neq k$ and on any extra measurement outcome $o_{\text{extra}}$ obtained by Alice, for all $k\in[N]$. Crucially, we note that the result applies to arbitrary quantum strategies by Alice, including arbitrary joint quantum measurements on the quantum states received in all $N$ rounds.
@@ 2 Estimation of experimental imperfections
where $N_{tu}$ is the total number of pulses prepared by Bob in the state labelled by $t$ in the basis labelled by $u$ that were measured by Alice in the basis labelled by $u$, and where $N_{tu}^{\text{error}}$ is the number of Alice’s incorrect measurement outcomes $t\oplus 1$ obtained from the $N_{tu}$ pulses, for all $t,u\in\{0,1\}$. Given that the frequencies $\bar{E}_{tu}$ are obtained from binomial distributions, their standard deviations are
@@
Table S1: Statistics for the average error rates.
@@ 2.5 Derivation of an upper bound on $\mu$ from a lower bound on $\eta_{\text{B}}$
3 Upper bound on the uncertainty angle in the preparation state
In this section we describe the experimental procedure used to determine an upper bound $\theta$ for the uncertainty angle on the Bloch sphere for Bob’s prepared states, and an upper bound $P_{\theta}$ on the probability that this bound is not satisfied.
As shown in Fig. 1 of the main text, Bob prepared the quantum states using two Pockels cells to modulate the bases and encoded bits. The Pockels cells were driven by quantum random number generators (QRNGs). We label the four target states by ‘0’, ‘1’, ‘+’ and ‘-’. Alice used a half-wave plate (HWP) and a polarizing beam splitter (PBS) to measure the quantum states. We used Alice’s setup to measure the quantum states in one of two bases by setting the HWP at one of two possible angles using a rotation mount. We experimentally estimated $\theta$ and $P_{\theta}$ using this joint setup, with some variations discussed below. Given the experimental setup, we refer to these below as measurements by Alice. Note however that in a real world implementation, these estimations would be performed by Bob using his own independent measurement setup.
To estimate $\theta$, Alice’s two single photon detectors were replaced by two power meters; and the intensity of the incoming light pulse was set to the higher value of approximately 18 mW. For each of the four states prepared by Bob, we measured the intensity of light measured by each of the two power meters. We repeated this 1000 times for each of the four states prepared by Bob. If Bob prepared the target states perfectly and the experimental setup were ideal then only one of the two power meters would measure a non-zero value. However, due to imperfections of the preparation procedure and the experimental setup, both power meters measure non-zero values, although one is much smaller than the other.
@@ 3 Upper bound on the uncertainty angle in the preparation state
to six decimal places. The uncertainty angle for Bob’s state preparation would be given by $\theta=\max_{i\in\{0,1,+,-\}}\{\alpha_{i}\}$, giving $\theta={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3.429185^{\circ}}$. However, we need to consider the imperfections of the PBS, HWP and rotation mount to derive a more accurate value of $\theta$, as discussed below.
3.1 Considering the experimental imperfections of the PBS, HWP and rotation mount
We now provide a derivation of $\theta$ and $P_{\theta}$ considering the experimental imperfections of the PBS, HWP and rotation mount using the experimental setup depicted in Fig. S1.
@@ 3.1 Considering the experimental imperfections of the PBS, HWP and rotation mount
We placed the PBS before the polarizer, as shown in Fig. S1(b). The light after the PBS is supposed to be horizontally(vertically)-polarized. We set the polarizer to angles of 0 and 90 degrees to test the actual polarization. We measured the residual vertical component when the light is supposed to be horizontally-polarized, and vice versa. We repeated this procedure ten times. We obtained a contrast of intensities given by
@@
Now let us consider that the states $\ket{+}$ or $\ket{-}$ are prepared. Thus, the HWP is aimed at $22.5^{\circ}$. In this case, if the HWP were perfect and perfectly aligned, the Bloch vector of the imperfectly prepared state would be rotated to the ‘attempted rotated state’ shown in Fig. S2(b). Due to the imperfections and misalignment of the HWP, the state is in fact rotated to the ‘rotated state’ illustrated in Fig. S2(b). Fig. S2(b) illustrates the worst case scenario that allows us to derive an upper bound on $\theta$. As in Fig. S2(a), we assume the two Bloch vectors lie on the real plane defined by the PBS axes. Thus, as above, (72) and (73) hold.
@@ 4 Time sequence and transaction times
Table S2 gives details of the obtained time measurements relevant for computing QA in our intracity implementation in Jinan, with laboratories communicated by $L_{\text{fibre}}=2,766$ m of optical fibre and physically separated by 425 m. The average value for the transaction time $\Delta T_{\text{tran}}$ was $15.34~{}\mu s$. The transaction time of the classical cross-checking scheme discussed in the Methods section is given by $\Delta T_{\text{tran,C}}=2\Delta T_{\text{comm}}$ (Eq. (8) in Methods), where $\Delta T_{\text{comm}}=\frac{L_{\text{fibre}}}{c_{\text{fibre}}}$ is the time that it takes to communicate a bit between $L_{0}$ and $L_{1}$ over the optical fibre channel, which took the value $\Delta T_{\text{tran,C}}=27.66~{}\mu s$ in our implementation in Jinan, and where $c_{\text{fibre}}=2\times 10^{8}$ ms-1 is the speed of light through the optical fibre. Thus, we obtained a quantum advantage of $QA=\Delta T_{\text{tran,C}}-\Delta T_{\text{tran}}=27.66~{}\mu\text{s}-15.34~{}\mu\text{s}=12.32~{}\mu\text{s}$ in our intracity implementation in Jinan.
Table S3 gives details of the obtained time measurements relevant for computing CA in our intercity implemetation between Yiyuan and Mazhan, with laboratories communicated by 60.54 km field-deployable optical fibre and physically separated by $D=51.60$ km. The average value for the transaction time $\Delta T_{\text{tran}}$ was $304.20~{}\mu$s. The transaction time of the classical cross-checking scheme discussed in the Methods section is given by $\Delta T_{\text{tran,CF}}=\frac{2D}{c}=344~{}\mu$s (Eq. (9) in Methods), where $c=3\times 10^{8}~{}$ms-1 is the speed of light through a vacuum. Thus, we obtained a comparative advantage of $CA=\Delta T_{\text{tran,CF}}-\Delta T_{\text{tran}}=344~{}\mu\text{s}-304.20~{}\mu\text{s}=~{}39.80~{}\mu\text{s}$ in our intercity implementation.
@@ 5 Related work and extensions of our schemes
Bartkiewicz et al. [24] describe an experimental implementation of partial cloning attacks on photon states representing components of quantum money tokens. The tokens were very short-lived, as they did not implement quantum memory. Bozzio et al. [26] demonstrated an on-the-fly version of quantum money tokens using weak coherent states of light. Guan et al. [27] also implemented short-lived quantum money with light pulses, using high-dimensional time-bin qudits. Again, neither of these implementations were integrated with quantum memory, and so the tokens were very short-lived. Jirakova et al. [28] presented proof-of-concept attacks on the implementation reported in Ref. [26].
@@
Schiansky et al. [38] recently reported an experimental demonstration of quantum digital payments using an interesting scheme that uses some features of quantum S-money. Ref. [38] also offers comments on and comparisons with our scheme and a previous partial implementation [14]. It is important to note that the proposed use case of the scheme of Ref. [38] is very different from that on which we focus here, and its suggested advantages are intended to apply in that use case. We comment briefly here only on some key points relevant to our present discussion, leaving a more complete comparison and analysis of the two schemes in various use cases for future work.
The scheme of Ref. [38] is not described in a relativistic context and not designed for the scenarios discussed in the present work. It involves three types of party, banks (trusted token providers or TTPs in the terminology of Ref. [38]), clients and merchants. To compare directly to the scenarios and use cases discussed in the present work, it is helpful to simplify by eliminating the merchants (or alternatively considering them as untrusted classical channels). In that case, one or more TTPs would supply streams of quantum states over unauthenticated channels that are accessible to clients, who may carry out appropriate measurements to generate classical tokens.
If there are two or more spatially separated TTPs this allows a single client (with multiple agents) to generate independent classical tokens from each stream and present each to the relevant TTP, who will validate them without cross-checking. The client can thus fraudulently multiply spend a single credit. Hence the scheme does not protect against multiple spending in the scenarios considered in the present work. (Its aim is to give a different type of protection, namely to prevent multiple transactions with the same identifier/label, without the need to cross-check.) Protection could be obtained by cross checking, but in that case there is no time advantage compared to a purely classical cross checking scheme. Alternatively, unforgeability can be guaranteed if there is a single TTP, as in the experimental demonstration of Ref. [38]. But in this case, two-way communication is required between the client and the TTP. Again this gives no time advantage over classical cross-checking schemes.
@@
The scheme of Ref. [38] requires clients to first decide where to spend their credit and then generate classical token data by carrying out appropriate measurements on a string of quantum states. In contrast, our recent [14] and current implementation allow the user flexibility to make this choice shortly before presenting the token for verification. The time between choice and presentation is limited in principle only by causal communication constraints, and in practice we achieved the sub-millisecond transaction times reported above. Our implementation allows the quantum communication between bank and user to take place arbitrarily long (even years) before the user chooses when and where to spend the token. In this respect it replicates the real-world functionality of money and credit cards, which are typically obtained from banks without any need to commit to using them with a given merchant or at a given time and place.
The implementation of Ref. [38] required “a few tens of minutes” [38] from Alice’s (the client’s) choice of presentation point to token validation. This should be compared with the $\approx 1.5~{}\mu{\rm s}$ required to validate a presented token in our experiments, the total transaction times of $\approx 15~{}\mu{\rm s}$ and $\approx 30{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}~{}\mu{\rm s}$ for our intra-city and inter-city experiments and also, importantly, with the times required for classical cross-checking schemes to complete a token transaction, which would be $\approx 28~{}\mu{\rm s}$ in our intracity setup using our 2,766 m long optical fibre link and $\approx 344~{}\mu{\rm s}$ in our intercity scenario using ideal light speed communication through free space with nodes separated by 51.651.6 51.651.6 km, and which would be $\approx 7~{}\mu{\rm s}$ for the 641 m long optical fibre link used in the experiment of Ref. [38] – or even shorter if the classical cross-checking scheme used ideal light speed communication through free space. Ref.  [38] suggests that with sufficiently good technology their token generation time could be reduced to the order of one second. Our transaction times are still far shorter than this target, which is also considerably longer than the classical cross-checking time required even for a global network. For comparison, great circle free space communication between antipodal Earth points would take about 67 ms.
It may be challenging to extend the scheme of Ref. [38] to networks with multiple transaction nodes and longer distances between nodes. One major challenge is overcoming quantum channel losses. The experiment of Ref. [38] implemented quantum communication through a 641 metres optical fibre with a total loss of $\approx 22\%$, of which $\approx 7\%$ corresponded to the fibre. With their reported error rate of $\approx 3.3\%$, their results are just within the secure region for the simplest case of two merchants, as illustrated in Fig. 4 of Ref. [38]. For realistic use cases with multiple merchants and large scale networks, higher losses would make it harder to obtain security. In contrast, quantum channel losses do not scale with network size in our implemented protocol, as all quantum communications may take place long before token use, over a single short dedicated channel with a known low loss rate. The communications between distant nodes in our protocol are purely classical. We analyze and discuss in section 7 below how our implementation can be extrapolated to multi-node global networks and still achieve good security.